# Adversarial Training for Multimodal Large Language Models against Jailbreak Attacks

### v1

@@ 1 Introduction
To address these challenges, we propose Projection Layer Against Adversarial Training (ProEAT), an innovative end-to-end adversarial training framework 2 designed to enhance the robustness of MLLMs against jailbreak attacks. ProEAT first incorporates adversarial training focused on a lightweight visual projection module, optimizing its parameters through a mean square error (MSE) objective by processing both original and adversarial samples, significantly reducing computational overhead while maintaining effectiveness. Additionally, we introduce a dynamic weight adjustment mechanism to dynamically balance loss function components, further streamlining parameter tuning and improving training efficiency. To enhance performance, ProEAT adopts a joint optimization strategy 3 across visual and textual modalities, ensuring cross-modal consistency and robust resistance to adversarial attacks originating from either modality. This comprehensive framework effectively addresses both efficiency and performance challenges, enabling secure and robust multimodal feature extraction and response generation.
@@
  • We propose an innovative end-to-end joint training framework that integrates projector and language model components, enabling collaborative optimization across modalities and significantly improving defense performance.
@@ 2.1 Multimodal Large Language Models
Recent advances have introduced more innovative fusion techniques. For instance, LLaMA-Adapter V2 [15] achieves cross-modal interaction through lightweight adaptation prompts, offering improved flexibility without significant computational overhead. CogVLM [45] takes a more intensive approach by incorporating visual expert modules directly into the attention and feedforward network layers, allowing deeper integration of visual and textual features. Although these MLLMs have demonstrated impressive performance across various tasks, their increasing deployment in critical applications has raised important security concerns, particularly regarding their vulnerability to adversarial attacks and cross-modal manipulations.
@@ 2.3 Adversarial Training for MLLMs
Adversarial training has become one of the most effective methods to defend against escape attacks from MLLMs. Early training-based defense strategies mainly improve the defense capability of models through parameter fine-tuning. For example,  [22, 6] first attempted supervised fine-tuning by mixing harmful and harmless prompts, but this method is difficult to defend against automated attacks. Subsequent studies have overcome these limitations by improving attack prompts [11], using gradient ascent with affirmative responses [5], and eliminating harmful knowledge [20, 56]. However, existing adversarial training methods still face challenges in dealing with the gap between continuous representations and discrete texts.
To address these challenges, we propose a novel end-to-end joint training framework to enhance MLLMs’ robustness against jailbreak attacks. Our approach first performs adversarial training on the projector by processing both original and noise-perturbed samples, optimizing through MSE minimization. We then innovatively integrate the projector and language model into a unified feature space, enabling robust multi-modal representations that resist adversarial perturbations. This joint training framework, combined with dynamic loss weighting, significantly improves the model’s defensive capabilities through cross-modal consistency and feature-level supervision.

### v2

@@ 1 Introduction
To address these challenges, we propose Projection Layer Against Adversarial Training (ProEAT), an innovative end-to-end adversarial training framework 2 designed to enhance the robustness of MLLMs against jailbreak attacks. ProEAT first incorporates adversarial training focused on a lightweight visual projection module, optimizing its parameters through a mean square error (MSE) objective by processing both original and adversarial samples, significantly reducing computational overhead while maintaining effectiveness. Additionally, we introduce a dynamic weight adjustment mechanism to dynamically balance loss function components, further streamlining parameter tuning and improving training efficiency. To enhance performance, ProEAT adopts a joint optimization strategy 3 across visual and textual modalities, ensuring cross-modal consistency and robust resistance to adversarial attacks [28, 29, 59, 30, 26] originating from either modality. This comprehensive framework effectively addresses both efficiency and performance challenges, enabling secure and robust multimodal feature extraction and response generation.
@@
  • We present a novel unified framework that jointly trains projector and language model modules through cross-modal co-optimization, substantially enhancing defensive efficacy against adversarial attacks.
@@ 2.1 Multimodal Large Language Models
Recent advances have introduced more innovative fusion techniques. For instance, LLaMA-Adapter V2 [15] achieves cross-modal interaction through lightweight adaptation prompts, offering improved flexibility without significant computational overhead. CogVLM [55] takes a more intensive approach by incorporating visual expert modules directly into the attention and feedforward network layers, allowing deeper integration of visual and textual features. Although these MLLMs have demonstrated impressive performance across various tasks, their increasing deployment in critical applications has raised important security concerns [31, 27, 67], particularly regarding their vulnerability to adversarial attacks and cross-modal manipulations.
@@ 2.3 Adversarial Training for MLLMs
Adversarial training [32, 33, 69, 52, 34, 25] has become one of the most effective methods to defend against escape attacks from MLLMs. Early training-based defense strategies mainly improve the defense capability of models through parameter fine-tuning. For example,  [22, 6] first attempted supervised fine-tuning by mixing harmful and harmless prompts, but this method is difficult to defend against automated attacks. Subsequent studies have overcome these limitations by improving attack prompts [11], using gradient ascent with affirmative responses [5], and eliminating harmful knowledge [20, 70]. However, existing adversarial training methods still face challenges in dealing with the gap between continuous representations and discrete texts.
To address these challenges, we propose ProEAT, an end-to-end collaborative training framework designed to enhance the jailbreak attack defense capabilities of MLLMs. The methodology initiates with adversarial projector training, where dual-stream processing of pristine and adversarially perturbed inputs optimizes modality alignment via MSE-driven gradient descent. Subsequently, we establish parameter-shared coupling between the projector and language model in a unified embedding space, enabling robust multi-modal representations that resist adversarial perturbations. This joint training framework employs dynamic multi-task optimization that enforces cross-modal semantic consistency constraints while providing feature-level supervision of latent space geometry and attention patterns. Through this joint adversarial training approach with adaptive loss weighting, ProEAT achieves an average of 34% stronger defense efficacy against jailbreak attacks compared to conventional methods.